# I Know About “Up”! Enhancing Spatial Reasoning in Visual Language Models Through 3D Reconstruction

### v1

@@ Abstract
I Know About “Up”! Enhancing Spatial Reasoning in Visual Language Models Through 3D Reconstruction
Hao Zhou Guangdong Polytechnic Normal University hao.zhou28@outlook.com                        Zaiqiao Mengthanks: Corresponding Author. University of Glasgow zaiqiao.meng@glasgow.ac.uk                        Yifang Chen Guangdong Polytechnic Normal University chenyf@gpnu.edu.cn
1 Introduction
Visual Language Models (VLMs), such as LLaVA (Liu et al., 2024), MiniGPT-4 (Zhu et al., 2023) and InternGPT (Liu et al., 2023d), are a class of deep neural models adept at simultaneously processing and understanding both visual and linguistic information. These VLMs often consist of an image encoder, an embedding projector to align image and text representations and a text decoder to process the projected image embedding and text representations, enabling joint understanding and reasoning between these modalities (Li et al., 2022). By using the language generation power of underlining text decoders, these VLMs showcase remarkable interaction capabilities in various applications, such as referring expression comprehension (Li et al., 2019b; Yu et al., 2016), visual question answering (Goyal et al., 2017; Zellers et al., 2019; Hudson and Manning, 2019), visual language reasoning (Liu et al., 2022a), and entailment (Liu et al., 2021; Xie et al., 2019).
Figure 1: An example of the VQA task, where humans can easily recognize positions under different views, but the vanilla LLaVA (Liu et al., 2024) can only predict correctly from certain views. By performing 3D reconstruction to obtain different views of the image, we can improve LLaVA’s predictive accuracy.
@@ 1 Introduction
Presently, numerous studies explore various benchmarks and methods to enhance the visual spatial reasoning capabilities of VLMs to improve their performance, exemplified by visual spatial reasoning (Liu et al., 2023a) and Whatsup_vlm (Kamath et al., 2023). These endeavors employ diverse methodologies (Su et al., 2019; Qi et al., 2020; Kiela et al., 2019) to augment the visual spatial reasoning abilities of models. However, despite these efforts yielding improvements in model performance to some extent, substantial challenges persist in comprehending complex scenes and intricate spatial relationships. Existing methodologies predominantly rely on 2D information, impeding the comprehensive capture of 3D spatial relationships among objects. For example, Figure 1 illustrates an example of the VQA task. While humans could easily recognize positions under different views of the objects, LLaVA can only predict correctly from certain views due to the lack of image views during their pretraining.
To tackle this challenge, we propose a novel model, called ZeroVLM, to solve the visual spatial reasoning task through the 3D reconstruction technique, entailing the reconstruction of images in 3D and capturing them from various views. In particular, our ZeroVLM utilize the Zero-1-to-3 (Liu et al., 2023c) model for constructing 3D views from a single 2D image within our tested datasets. This allows VLMs to access richer spatial information through 3D transformation, thereby enhancing their visual spatial reasoning capabilities. We conducted comprehensive experiments across four visual spatial reasoning datasets, and the results show that all the tested VLMs have notably improved spatial reasoning capabilities through our ZeroVLM.
@@
Figure 2: Single-view images were generated using Zero-1-to-3 to produce left-view, right-view, and random-view images. Multi-view images were created by combining these different single-view images in various configurations.
2 Related Work
Vision Language Models (VLMs). VLMs amalgamate computer vision and natural language processing technologies to comprehend and generate correlations between images and natural language (Dosovitskiy et al., 2020). VLMs can accept image input to generate corresponding response in text (Menon and Vondrick, 2022). In terms of training architecture, recent works include joint pretraining architectures (e.g. OFA Wang et al. (2022) and VLMo (Bao et al., 2021)), which train image and text data jointly, and image-to-text mapping architectures (e.g., PaLI (Chen et al., 2022b)), which map an image encoder to a well-pretrained text encoder. The latter approach has gained more popularity. VLMs exhibit cross-modal understanding capabilities, enabling them to extract information from images and translate it into natural language, or retrieve information from natural language and generate corresponding images (Chen et al., 2022a). This wide range of capabilities broadens the potential of VLMs (e.g., GlaMM (Rasheed et al., 2023) and MiniGPT-4 (Zhu et al., 2023)) in understanding and generating associations between images and language, such as Image Captioning (Zhang et al., 2021), Visual Question Answering (Yang et al., 2022), and Multi-modal Translation (Li et al., 2019a). In this study, we aim to explore the visual-spatial recognition capabilities of these VLMs.
@@ 2 Related Work
3D Reconstruction. Visual spatial reasoning encompasses the cognitive ability to comprehend and manipulate visual spatial information, while 3D reconstruction involves generating a three-dimensional representation of an object or scene from two-dimensional images or other sensor data. For instance, MonoScene (Cao and De Charette, 2022) infers dense geometric structures and semantic information of a scene from a single monocular RGB image, offering an efficient and innovative approach to complete 3D semantic scenes. VoxFormer (Li et al., 2023) proposes a transformer-based (Vaswani et al., 2017) semantic scene completion framework, addressing camera-based 3D semantic scene completion by introducing sparse voxel queries (Guchait et al., 2021) and masked autoencoder design (He et al., 2022). Zero-1-to-3 (Liu et al., 2023c) is a 3D reconstruction model that excels in zero-shot generalization on out-of-distribution datasets and unprocessed images by using a single image as input without requiring additional 3D or depth information. We chose this model for its simplicity and effectiveness compared to MonoScene (Cao and De Charette, 2022) and VoxFormer (Li et al., 2023).
Figure 3: An overview of our proposed ZeroVLM model. Our ZeroVLM first uses Zero-1-to-3 to perform 3D reconstruction to obtain different views of the input image, and then it stitches the original images with these different views to obtain the stitched image, which is the input of a VLM for answer prediction.
3 Preliminaries
In this section, we will formally define the task of visual spatial reasoning and introduce the Zero-1-to-3 model (Liu et al., 2023c), which is one of the components utilized by our model.
3.1 Visual Spatial Reasoning Task
@@ 3.2 Zero-1-to-3
Zero-1-to-3 (Liu et al., 2023c) is a model designed for visual reconstruction tasks. It works by identifying objects in input image i and adjusting the camera perspective of these objects. Given a image $i\in\mathbb{R}^{H\times W\times 3}$ and a relative camera rotation matrix $R\in\mathbb{R}^{3\times 3}$ and a translation vector $T\in\mathbb{R}^{3}$, Zero-1-to-3 leans a model $f$ such that its output is a new perspective image under the specified camera transformation:
$\hat{i}_{R,T}=f(i,R,T)$ (1)
@@
where $\hat{i}_{R,T}$ represents the generated new perspective image. When we input original image i and hope to generate a 3D reconstructed view from the left at an angle of 45° through Zero-1-to-3, we can achieve this effect by setting $R$ to the rotation matrix of 45° around the Y axis and setting $T$ to the vector translated to the left by a certain distance. In order to synthesize new views under such partial constraints, Zero-1-to-3 uses a large-scale diffusion model to integrate the geometric prior knowledge of the input image. The conditional diffusion model embedded in Zero-1-to-3  (Yang et al., 2023; Podell et al., 2023) learns to control the relative camera viewpoint using synthetic datasets, which helps generate new images with specified camera transformations, such as fixed camera viewpoints (such as left or right view) or randomly generated camera viewpoints. Despite being trained on synthetic datasets, Zero-1-to-3 still demonstrates strong zero-shot generalization capabilities to out-of-distribution datasets and real-world images. Figure 2 shows the different views generated by Zero-1-to-3 based on the input image.
4 Methodologies
4.1 Overview of ZeroVLM
@@
To address the visual spatial reasoning task, we propose ZeroVLM, a model that leverages Zero-1-to-3 to infer various views of input images and employs a VLM to generate answers using the combined multiview images and our specially designed view prompts. The overview of our model is illustrated in Figure 3. In particular, ZeroVLM is a novel visual-language model that combines large language models (such as LLaMA (Touvron et al., 2023)) with high-level vision models, aiming to enhance visual-spatial reasoning capabilities by leveraging 3D reconstruction techniques. Given an image i, ZeroVLM uses Zero-1-to-3 to infer its various views $\hat{i}_{R,T}$ and performs image stitching from multiple views, thereby providing richer spatial information. The 3D reconstruction feature provided by Zero-1-to-3 can enable ZeroVLM to better understand and infer spatial relationships by viewing objects from different angles. At the same time, a specialized view prompt is designed into ZeroVLM to further enhance its visual spatial reasoning capabilities. This prompt helps guide the model to focus on relevant spatial views and improve its interpretation accuracy. The architecture of our ZeroVLM is shown in Figure 3.
4.2 Data Augmentation by 3D Reconstruction
@@
In our work, we use Zero-1-to-3 for 3D reconstruction on the dataset to generate $\hat{i}_{R,T}$ from different viewpoints, e.g. the left, right and random views. Our investigation not only explores whether the creation of single-view images can enhance the visual spatial reasoning capabilities of VLMs but also examines whether multi-view images can help this improvement, where multi-view images are synthesized from different $\hat{i}_{R,T}$. Because multi-view images can provide richer spatial information, the model can get the opportunity to observe the same scene from different angles, thereby capturing a more comprehensive spatial layout and the relationship between objects, and different views can provide redundant information, so that the model can still make accurate judgments when facing noise or partial information loss. Therefore, in our work, after inferring different views of the input images, we further construct multi-view images from this single views by stitching them, and test their effectiveness against the single view images. Figure 2 shows an example of the multi-view image constructed by stitching multiple single view images generated through Zero-1-to-3.
@@
By synthesizing these various multi-view images, we aim to determine their effectiveness in improving the spatial reasoning abilities of VLMs.
4.3 View Prompt
@@
In our work, we first test whether 3D reconstruction of images can improve the accuracy of VLM’s visual spatial ability at the image level. To further explore whether the context prompt can enhance VLM’s visual spatial reasoning ability, we introduced a special prompt called view prompt in the experiment. This view prompt varies depending on the input image and its views. Figure 4 shows two view prompt examples over a single view and multiple views of the input.
@@
We designed a variety of view prompts based on the content of different view images to guide VLM to better understand and reason about the view spatial relationship between the target object and other objects in the image. We first used Zero-1-to-3 (Liu et al., 2023c) to perform 3D reconstruction of VSR (Liu et al., 2023a) dataset and What’sUp (Kamath et al., 2023) dataset from different viewpoints. Then during the inference of VLM, we use a prompt consisting of the question, the view prompt, and the stitched view image, to generate the answer. Figure 5 provides a detailed description of the process.
@@
Figure 4: These view prompts are manually constructed by us. View prompt comparison between single-view images and multi-view images. {question} is the corresponding question in the prompt.
@@
Figure 5: For both single-view and multi-view datasets, we employed a view prompt. The aim was to explore whether perspective language models could enhance their visual spatial reasoning abilities through improvements at the textual level.
5 Experiments
5.1 Experimental Setup
@@
Table 1: The number of images in various datasets.
Baselines. In our research, we employ LLaVA and MiniGPT-4 as the primary VLMs. LLaVA utilizes multi-modal language-image instruction data for instruction adjustment, employing CLIP-ViT-L-336px (Liu et al., 2023b) as the visual encoder and MLP projection as the visual language cross-modal connector, achieving comprehensive understanding of both visual and language inputs. The architecture of MiniGPT-4 involves aligning a frozen visual encoder with a frozen high-level language model, e.g. Vicuna (Chiang et al., 2023), through projection layers, ensuring correct alignment of visual features with the high-level large language model. MiniGPT-4 can exhibit advanced multi-modal capabilities similar to those of GPT-4 (Achiam et al., 2023).
@@
Human 95.4 95.4 100 100
CLIP (frozen) 56.0 54.50 - -
CLIP (FT) 65.10 - - -
VisualBERT 55.20 51.00 - -
ViLT 69.30 63.00 - -
LXMERT 70.10 61.20 - -
@@
Table 2: Accuracy performance of the compared models. Missing values indicate that ZeroVLM (L) and ZeroVLM (M) were not tested on the visual spatial reasoning on the VSR and What’sUp datasets. Bests are in bold, and second bests are underlined.
@@
Figure 6: L-V represents the left-view images, R-V represents the right-view images, Ra-V represents the random-view images, and M-V represents the multi-view images (excluding the original image). Note that here we use ZeroVLM (L) and ZeroVLM (M) for fair comparison.
5.2 Experimental Result
Overall performance. Our experimental aims to comprehensively evaluate the visual spatial reasoning capabilities of VLMs ZeroVLM (L) and ZeroVLM (M) on various datasets. We seek to understand how different datasets (including original, single-view, and multi-view datasets) affect the performance of these models in visual-spatial reasoning tasks. We first test ZeroVLM (L) and ZeroVLM (M) on the original dataset (without 3D reconstruction) and compare the results with those of VSR (Liu et al., 2023a). The result of this comparison is reported in Table 2. Our experimental results show a baseline comparison of ZeroVLM (L) and ZeroVLM (M) testing the original dataset with VSR (Liu et al., 2023a), highlighting the basic performance of ZeroVLM (L) and ZeroVLM (M).

### v2

@@ Abstract.
Visual Spatial Reasoning, 3D Reconstruction, Visual Language Models
isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Visual Language Models (VLMs), such as LLaVA (Liu et al., 2024), MiniGPT-4 (Zhu et al., 2023) and InternGPT (Liu et al., 2023b), are a class of deep neural models adept at simultaneously processing and understanding both visual and linguistic information. These VLMs often consist of an image encoder, an embedding projector to align image and text representations and a text decoder to process the projected image embedding and text representations, enabling joint understanding and reasoning between these modalities (Li et al., 2022). By using the language generation power of underlining text decoders, these VLMs showcase remarkable interaction capabilities in various applications, such as referring expression comprehension (Li et al., 2019b; Yu et al., 2016), visual question answering (Goyal et al., 2017; Zellers et al., 2019; Hudson and Manning, 2019), visual language reasoning (Liu et al., 2022a), and entailment (Liu et al., 2021; Xie et al., 2019).
Figure 1. An example of the VQA task, where humans can easily recognize positions under different views, but the vanilla LLaVA (Liu et al., 2024) can only predict correctly from certain views. By performing 3D reconstruction to obtain different views of the image, we can improve LLaVA’s predictive accuracy.
@@ 1. Introduction
Presently, numerous studies explore various benchmarks and methods to enhance the visual spatial reasoning capabilities of VLMs to improve their performance, exemplified by visual spatial reasoning (Liu et al., 2023a), Whatsup_vlm (Kamath et al., 2023) and SpatialVLM (Chen et al., 2024). These endeavors employ diverse methodologies (Su et al., 2019; Qi et al., 2020; Kiela et al., 2019) to augment the visual spatial reasoning abilities of models. However, despite these efforts yielding improvements in model performance to some extent, substantial challenges persist in comprehending complex scenes and intricate spatial relationships. Existing methodologies predominantly rely on 2D information, impeding the comprehensive capture of 3D spatial relationships among objects. For example, Figure 1 illustrates an example of the VQA task. While humans could easily recognize positions under different views of the objects, LLaVA can only predict correctly from certain views due to the lack of image views during their pretraining.
To tackle this challenge, we propose a novel model, called ZeroVLM, to solve the visual spatial reasoning task through the 3D reconstruction technique, entailing the reconstruction of images in 3D and capturing them from various views. In particular, our ZeroVLM utilize the Zero-1-to-3 (Liu et al., 2023d) model for constructing 3D views from a single 2D image within our tested datasets. This allows VLMs to access richer spatial information through 3D transformation, thereby enhancing their visual spatial reasoning capabilities. We conducted comprehensive experiments across four visual spatial reasoning datasets, and the results show that all the tested VLMs have notably improved spatial reasoning capabilities through our ZeroVLM.
@@
Figure 2. Single-view images were generated using Zero-1-to-3 to produce left-view, right-view, and random-view images. Multi-view images were created by combining these different single-view images in various configurations.
2. Related Work
Vision Language Models (VLMs). VLMs amalgamate computer vision and natural language processing technologies to comprehend and generate correlations between images and natural language (Dosovitskiy et al., 2020). VLMs can accept image input to generate a corresponding response in text (Menon and Vondrick, 2022). In terms of training architecture, recent works include joint pretraining architectures (e.g. OFA (Wang et al., 2022) and VLMo (Bao et al., 2021)), which train image and text data jointly, and image-to-text mapping architectures (e.g., PaLI (Chen et al., 2022b)), which map an image encoder to a well-pretrained text encoder. The latter approach has gained more popularity. VLMs exhibit cross-modal understanding capabilities, enabling them to extract information from images and translate it into natural language, or retrieve information from natural language and generate corresponding images (Chen et al., 2022a). This wide range of capabilities broadens the potential of VLMs (e.g., GlaMM (Rasheed et al., 2023) and MiniGPT-4 (Zhu et al., 2023)) in understanding and generating associations between images and language, such as Image Captioning (Zhang et al., 2021), Visual Question Answering (Yang et al., 2022), and Multi-modal Translation (Li et al., 2019a). In this study, we aim to explore the visual-spatial recognition capabilities of these VLMs.
@@ 2. Related Work
3D Reconstruction. Visual spatial reasoning encompasses the cognitive ability to comprehend and manipulate visual spatial information, while 3D reconstruction involves generating a three-dimensional representation of an object or scene from two-dimensional images or other sensor data. For instance, MonoScene (Cao and De Charette, 2022) infers dense geometric structures and semantic information of a scene from a single monocular RGB image, offering an efficient and innovative approach to complete 3D semantic scenes. VoxFormer (Li et al., 2023) proposes a transformer-based (Vaswani et al., 2017) semantic scene completion framework, addressing camera-based 3D semantic scene completion by introducing sparse voxel queries (Guchait et al., 2021) and masked autoencoder design (He et al., 2022). Zero-1-to-3 (Liu et al., 2023d) is a 3D reconstruction model that excels in zero-shot generalization on out-of-distribution datasets and unprocessed images by using a single image as input without requiring additional 3D or depth information. We chose this model for its simplicity and effectiveness compared to MonoScene (Cao and De Charette, 2022) and VoxFormer (Li et al., 2023).
Figure 3. An overview of our proposed ZeroVLM model. Our ZeroVLM first uses Zero-1-to-3 to perform 3D reconstruction to obtain different views of the input image, and then it stitches the original images with these different views to obtain the stitched image, which is the input of a VLM for answer prediction.
3. Preliminaries
In this section, we will formally define the task of visual spatial reasoning and introduce the Zero-1-to-3 model (Liu et al., 2023d), which is one of the components utilized by our model.
3.1. Visual Spatial Reasoning Task
@@ 3.2. Zero-1-to-3
Zero-1-to-3 (Liu et al., 2023d) is a model designed for visual reconstruction tasks. It works by identifying objects in input image i and adjusting the camera perspective of these objects. Given an image $i\in\mathbb{R}^{H\times W\times 3}$ and a relative camera rotation matrix $R\in\mathbb{R}^{3\times 3}$ and a translation vector $T\in\mathbb{R}^{3}$, Zero-1-to-3 leans a model $f$ such that its output is a new perspective image under the specified camera transformation:
(1) $\hat{i}_{R,T}=f(i,R,T)$
@@
where $\hat{i}_{R,T}$ represents the generated new perspective image. When we input the original image i and hope to generate a 3D reconstructed view from the left at an angle of 45° through Zero-1-to-3, we can achieve this effect by setting $R$ to the rotation matrix of 45° around the Y axis and setting $T$ to the vector translated to the left by a certain distance. In order to synthesize new views under such partial constraints, Zero-1-to-3 uses a large-scale diffusion model to integrate the geometric prior knowledge of the input image. The conditional diffusion model embedded in Zero-1-to-3  (Yang et al., 2023; Podell et al., 2023) learns to control the relative camera viewpoint using synthetic datasets, which helps generate new images with specified camera transformations, such as fixed camera viewpoints (such as left or right view) or randomly generated camera viewpoints. Despite being trained on synthetic datasets, Zero-1-to-3 still demonstrates strong zero-shot generalization capabilities to out-of-distribution datasets and real-world images. Figure 2 shows the different views generated by Zero-1-to-3 based on the input image.
4. Methodologies
4.1. Overview of ZeroVLM
@@
To address the visual spatial reasoning task, we propose ZeroVLM, a model that leverages Zero-1-to-3 to infer various views of input images and employs a VLM to generate answers using the combined multiview images and our specially designed view prompts. The overview of our model is illustrated in Figure 3. In particular, ZeroVLM is a novel visual-language model that combines large language models (such as LLaMA (Touvron et al., 2023)) with high-level vision models, aiming to enhance visual-spatial reasoning capabilities by leveraging 3D reconstruction techniques. Given an image i, ZeroVLM uses Zero-1-to-3 to infer its various views $\hat{i}_{R,T}$ and performs image stitching from multiple views, thereby providing richer spatial information. The 3D reconstruction feature provided by Zero-1-to-3 can enable ZeroVLM to better understand and infer spatial relationships by viewing objects from different angles. At the same time, a specialized view prompt is designed into ZeroVLM to further enhance its visual spatial reasoning capabilities. This prompt helps guide the model in focusing on relevant spatial views and improving its accuracy in interpretation. The architecture of our ZeroVLM is shown in Figure 3, with a detailed description of its components described in the following subsections.
4.2. Data Augmentation by 3D Reconstruction
@@
In our work, we use Zero-1-to-3 for 3D reconstruction on the dataset to generate $\hat{i}_{R,T}$ from different viewpoints, e.g. the left, right and random views. Our investigation not only explores whether the creation of single-view images can enhance the visual spatial reasoning capabilities of VLMs but also examines whether multi-view images can help this improvement, where multi-view images are synthesized from different $\hat{i}_{R,T}$. Because multi-view images can provide richer spatial information, the model can get the opportunity to observe the same scene from different angles, thereby capturing a more comprehensive spatial layout and the relationship between objects, and different views can provide redundant information, so that the model can still make accurate judgments when facing noise or partial information loss. Therefore, in our work, after inferring different views of the input images, we further construct multi-view images from these single views by stitching them and testing their effectiveness against the single-view images. Figure 2 shows an example of the multi-view image constructed by stitching multiple single-view images generated through Zero-1-to-3.
@@
By synthesizing these various single-view and multi-view images, we aim to conduct comprehensive controllable experiments to determine their effectiveness in improving the spatial reasoning abilities of VLMs.
4.3. View Prompt
@@
In our work, we first test whether the 3D reconstruction of images can improve the accuracy of VLM’s visual spatial ability at the image level. To further explore whether the context prompt can enhance VLM’s visual spatial reasoning ability, we introduced a special prompt called view prompt in the experiment. This view prompt varies depending on the input image and its views. Figure 4 shows two view prompt examples over a single view and multiple views of the input.
@@
We designed a variety of view prompts based on the content of different view images to guide VLM to better understand and reason about the view spatial relationship between the target object and other objects in the image. We first used Zero-1-to-3 (Liu et al., 2023d) to perform 3D reconstruction of VSR (Liu et al., 2023a) dataset and What’sUp (Kamath et al., 2023) dataset from different viewpoints. Then during the inference of VLM, we use a prompt consisting of the question, the view prompt, and the stitched view image, to generate the answer. Figure 5 provides detailed examples of the process.
@@
Figure 4. These view prompts are manually constructed by us. View prompt comparison between single-view images and multi-view images. {question} is the corresponding question in the prompt.
@@
Figure 5. For both single-view and multi-view datasets, we employed a view prompt. The aim was to explore whether perspective language models could enhance their visual spatial reasoning abilities through improvements at the textual level.
5. Experiments
5.1. Experimental Setup
@@
Table 1. The number of images in various datasets.
Baselines. In our research, we employ LLaVA and MiniGPT-4 as the primary VLMs. The LLaVA version we use is LLaVA-v1.5-13B and the MiniGpt-4 version is MiniGPT-4 (Vicuna 13B). LLaVA utilizes multi-modal language-image instruction data for instruction adjustment, employing CLIP-ViT-L-336px (Liu et al., 2023c) as the visual encoder and MLP projection as the visual language cross-modal connector, achieving comprehensive understanding of both visual and language inputs. The architecture of MiniGPT-4 involves aligning a frozen visual encoder with a frozen high-level language model, e.g. Vicuna (Chiang et al., 2023), through projection layers, ensuring correct alignment of visual features with the high-level large language model. MiniGPT-4 can exhibit advanced multi-modal capabilities similar to those of GPT-4 (Achiam et al., 2023).
@@
Human 95.40 95.40 100 100
CLIP (frozen) 56.00 54.50 58.00 57.50
CLIP (FT) 65.10 63.30 61.20 59.80
VisualBERT 55.20 51.00 53.40 55.60
ViLT 69.30 63.00 60.20 63.60
LXMERT 70.10 61.20 58.30 54.70
@@
Table 2. Accuracy performance of the compared models. Missing values indicate that ZeroVLM (L) and ZeroVLM (M) were not tested on the visual spatial reasoning on the VSR and What’sUp datasets. Bests are in bold, and second bests are underlined.
@@
Figure 6. L-V represents the left-view images, R-V represents the right-view images, Ra-V represents the random-view images, and M-V represents the multi-view images (excluding the original image). Note that here we use ZeroVLM (L) and ZeroVLM (M) for fair comparison.
5.2. Experimental Result
Overall performance. Our experimental aims to comprehensively evaluate the visual spatial reasoning capabilities of VLMs ZeroVLM (L) and ZeroVLM (M) on various datasets. We seek to understand how different view datasets (including original, single-view, and multi-view datasets) affect the performance of these models in visual-spatial reasoning tasks. We first test ZeroVLM (L) and ZeroVLM (M) on the original VSR dataset (without 3D reconstruction) and we reproduced LXMERT (Tan and Bansal, 2019), VisualBERT (Li et al., 2019b), and CLIP (Radford et al., 2021) using the same parameters as in VSR (Liu et al., 2023a), and also tested them on the original VSR dataset (without 3D reconstruction). The result of this comparison is reported in Table 2. Our experimental results show a baseline comparison of ZeroVLM (L) and ZeroVLM (M) testing the original dataset with VSR (Liu et al., 2023a), highlighting the basic performance of ZeroVLM (L) and ZeroVLM (M).